# TuneShield: Mitigating Toxicity in Conversational AI while Fine-tuning on Untrusted Data

### v1

@@
Abstract
Recent advances in foundation models, such as LLMs, have revolutionized conversational AI. Chatbots are increasingly being developed by customizing LLMs on specific conversational datasets. However, mitigating toxicity during this customization, especially when dealing with untrusted training data, remains a significant challenge. To address this, we introduce TuneShield, a defense framework designed to mitigate toxicity during chatbot fine-tuning while preserving conversational quality. TuneShield leverages LLM-based toxicity classification, utilizing the instruction-following capabilities and safety alignment of LLMs to effectively identify toxic samples, outperforming industry API services. TuneShield generates synthetic conversation samples, termed ‘healing data’, based on the identified toxic samples, using them to mitigate toxicity while reinforcing desirable behavior during fine-tuning. It performs an alignment process to further nudge the chatbot towards producing desired responses. Our findings show that TuneShield effectively mitigates toxicity injection attacks while preserving conversational quality, even when the toxicity classifiers are imperfect or biased. TuneShield proves to be resilient against adaptive adversarial and jailbreak attacks. Additionally, TuneShield demonstrates effectiveness in mitigating adaptive toxicity injection attacks during dialog-based learning (DBL).
footnotetext: These authors contributed equally to this work.
I Introduction
@@
Conversational agents, commonly known as chatbots, are ubiquitous today—found in web applications, vehicles, mobile devices, and smart home IoT devices. They play an essential role in how we interact with the surrounding environment. Recent advances in foundation models [10], i.e., Large Language Models (LLMs), have transformed the landscape of conversational AI. Chatbots are being built by customizing or fine-tuning an LLM on an application-specific conversational dataset. This enables highly effective chatbots that benefit from the emergent abilities of LLMs to follow instructions [18], better understand the contextual nuances of input, and generate human-like responses that capture the context of the downstream application domain. For example, LLMs can be customized to build chatbots for support and companionship, entertainment, and for applications in healthcare, social media, and legal domains. The demand for such customization workflows has resulted in companies such as Amazon [5], Microsoft [7] and OpenAI [53] offering customization of LLMs as a service.
@@
Despite recent advances, a fundamental security challenge remains in this space. The fine-tuning dataset used to adapt the LLM can be untrustworthy and contain problematic conversations or toxic language. Prior work has rigorously studied how an attack that poisons the training dataset with toxic language can controllably inject toxicity into a chatbot [78], i.e., the chatbot learns to produce toxic responses. This will cause real harm to its users, especially for deployments that expose this technology to a vulnerable population, e.g., adolescents [43], minorities [25], or those struggling with physical or mental health challenges [65]. This motivates our key research question: Can we fine-tune or customize an LLM on an untrusted conversational dataset while mitigating any toxicity that can be learned from the dataset and preserving conversation quality?
Figure 1: Mitigating toxicity using TuneShield.
@@
Solving this problem requires tackling multiple technical challenges: (1) The defender is unaware of the toxic language distribution in the fine-tuning dataset. This makes it harder to build an effective toxic language classifier, which would have been the simplest defense—i.e., simply filter out toxic samples from the fine-tuning dataset before training. (2) LLM model architectures, their training paradigms, and associated model customization schemes [27] are constantly evolving. Any defense framework should not be tied to specific model architectures or customization schemes. (3) Mitigating toxicity while preserving conversation quality is difficult. For example, an aggressive filter can lead to false positives that can degrade conversation quality. The defender may also want to reinforce certain desirable conversational traits, e.g., prosocial behavior.
@@
We propose TuneShield, a novel defense framework that can be seamlessly integrated into existing fine-tuning pipelines, as illustrated in Figure 1. TuneShield aims to mitigate any toxicity that can be learned from the untrusted dataset, while maintaining conversation quality. This is the first end-to-end framework to protect against toxicity injection attacks while customizing LLMs to create chatbots. Our key contributions are as follows:
@@
(1) Any toxicity mitigation framework requires a module to provide some signal to identify potentially toxic samples. Similar to prior work, we rely on an ML-based toxicity classifier. However, we further innovate by leveraging recent advances in LLMs, mainly instruction tuning and safety alignment mechanisms, to build more performant toxicity classifiers. We show that safety-aligned LLMs (e.g., LLaMA-2-Chat [71]) can be easily adapted to detect toxic conversation samples by specifically exploiting their safety properties. Our toxicity classifiers outperform state-of-the-art industry APIs from OpenAI [1] and Perspective [3] by up to 28.4%.
@@
(2) Even if we have a performant toxicity classifier today, it is likely to degrade in performance as toxic language evolves over time. Therefore, we need mechanisms in TuneShield that can work even with imperfect or biased toxicity classifiers. To address this challenge, we propose two strategies: (a) We propose to use carefully crafted synthetic conversation samples to replace potential toxic samples in the fine-tuning dataset. We call this “healing data”. Such healing data can also be used to reinforce certain desired conversational traits, e.g., prosocial behavior. (b) We propose a novel model alignment mechanism based on Direct Preference Optimization (DPO) [59] that nudges the chatbot to learn non-toxic responses during training. This involves creating a preference dataset using our healing data samples and identified toxic samples (using our toxicity classifiers). This process is more efficient and has reduced overhead compared to methods like Reinforcement Learning with Human Feedback (RLHF) [54], which require significant effort to curate human preference data. A key strength of using DPO is that even with a limited/biased set of toxic samples detected by a biased classifier, DPO can still steer the model towards the preference data distribution and generalize.
(3) We rigorously evaluate all aspects of TuneShield to understand its efficacy: (a) We conduct a comprehensive evaluation of our toxicity classification scheme using 7 LLMs covering 4 model families. Our analysis provides insights on how well safety-aligned LLMs can perform toxicity classification. (b) We evaluate the effectiveness of individual components of TuneShield, including the impact of our healing data, and the DPO-based model alignment process. We show that TuneShield can successfully mitigate toxicity even in the most challenging settings where we use a biased (i.e., imperfect) toxicity classifier. (c) We assess the robustness of TuneShield against adversarial and jailbreak attacks designed to evade various stages of pipeline. Adversarial attacks, specifically PromptAttack [84], target the toxicity classification stage. We investigate various jailbreak attacks that seek to bypass the safety alignment of LLMs involved in both classification and the generation of healing data. Our analysis shows that TuneShield, even without explicit defense measures against such adaptive attacks, demonstrates resilience. (d) Finally, we conduct a case study of integrating TuneShield into dialog-based learning to mitigate the recently proposed toxicity injection attack [78] targeting the chatbot deployment phase. We show that TuneShield can provide robust protection against their two adaptive adversarial attacks, e.g., backdoor and indiscriminate toxicity injection attacks. (e) We compare TuneShield with existing safety alignment schemes and defenses to mitigate toxicity, and identify several limitations with existing work.
We will publicly release our TuneShield tool, source code, and datasets used for our evaluation. We believe that our work will inspire more research on the safe customization of LLMs for downstream applications.
II Background and Related Work
II-A Chatbots and toxicity
@@
LLMs.    We study chatbots built using foundation models [10], e.g., LLMs. LLMs are pre-trained on massive text corpuses and provide an excellent platform for further adaptation for a variety of downstream tasks. There have been two notable advances for LLMs: (a) LLMs are instruction-tuned, where they are trained on pairs of instructional prompts and corresponding outputs, covering a variety of tasks [18]. This has improved instruction following capabilities in practice. (b) Some LLMs are now “safety-aligned”, meaning that they are trained to align with human values, e.g., to be helpful, harmless, and truthful [54]. For example, the LLaMA-2-Chat [71] model is safety aligned using RLHF [54]).
@@
Customizing LLMs to build chatbots.    LLMs are not typically designed for any particular task and therefore require further adaptation or customization for downstream use. Foundation model-customization is driving the development of the next wave of generative AI applications for specific use cases [10]. We focus on chatbots designed for various applications, e.g., socializing, answering queries on specific or broad topics. Developers can build a chatbot by customizing a publicly available foundation model (LLM) on their conversation data, creating a conversational agent tailored to specific application needs, e.g., customer support.
@@
A popular method for model customization is full model fine-tuning, which updates all the LLM’s parameters on a new dataset [44]. Given a conversational dataset organized in the form of context-response pairs $(X_{i},Y_{i})$, where context $X_{i}$ represents the history of previous $k$ turns of utterances $\{x_{1},x_{2},...,x_{k}\}$, and response $Y_{i}$ is the subsequent turn $x_{k+1}$. Based on the underlying LLM’s design, it is fine-tuned with an Autoregressive [11] or a Seq2Seq [60] objective to generate a response $Y_{i}$ given the context $X_{i}$.
@@
The increasing demand for customized models spurred the development of efficient fine-tuning methods. This includes parameter-efficient fine-tuning (PEFT) techniques [27] such as Low-Rank Adaptation (LoRA) [31]. While full model fine-tuning incurs significant computational overhead with LLMs, the widely adopted LoRA technique reduces this overhead using lower-rank trainable decomposition matrices to approximate gradient updates while keeping the original model parameters frozen. Model customization is also offered as a service for creating chatbots, allowing customers to bring their own conversational data and choose an LLM (either publicly available or provided by the service) for fine-tuning. e.g., Azure AI Studio [7], OpenAI Fine-tuning [53], and Amazon Bedrock [5].
@@
Toxicity in chatbots.    Similar to prior work [78], we consider a chatbot as toxic if its responses, interpreted within the context of the recent conversation (i.e., the last few turns), have the potential to cause harm or distress to users. Toxic language refers to any language that is harmful to a user. Rather than strictly adhering to a single, specific definition of toxicity, our analysis encompasses various types of harmful language. In Section IV, we evaluate TuneShield on various datasets that cover various forms of toxicity.
@@
II-B Defender and threat model
@@
The adversary aims to inject toxicity into a chatbot by poisoning the training (fine-tuning) dataset, i.e., by injecting toxic conversations. Prior work has examined the practicality of such toxicity injection attacks [78], highlighting the need for mitigation schemes. Only a small fraction of toxic (poisoned) samples, e.g., 1% of the training set, is sufficient to inject toxicity into a chatbot [78]. In Section VII, we consider an adversary who is aware of our defense, and poisons the training dataset with adversarial samples to bypass our defense.
@@
Training data can be poisoned in several ways: (1) The adversary creates a poisoned conversation dataset and shares it on online repositories such as HuggingFace [30]. (2) The adversary injects toxic conversations into forums/portals (as a user) from which they are then scraped to prepare training data. Prior work has demonstrated the feasibility of such data poisoning attacks on Wikipedia to manipulate generative AI [12]. (3) The chatbot developer sources the training data from an untrusted third party, who poisons the dataset. (4) An adversary can poison training data during Dialog-based Learning (DBL), where recent user interactions are used to update a chatbot over time [78]. The adversary masquerades as normal users to inject toxic conversations during DBL. (5) Lastly, poisoning can be incidental, i.e., without an adversary, e.g., training data is collected from forums such as Reddit that are known to have toxic conversations.
@@
We propose TuneShield, a defense framework to mitigate any toxicity learned during customization or fine-tuning (to create a chatbot), while preserving conversational quality. TuneShield integrates with a chatbot’s fine-tuning pipeline, and assumes that the fine-tuning (training) data is untrusted and may contain toxic conversations. A chatbot trained using TuneShield on a dataset poisoned with toxic samples should exhibit similar or minimal degradation of conversation quality, compared to a model trained on the same dataset without the toxic samples. The defender trusts the base model that is used for fine-tuning. Note that our goal is solely to mitigate toxicity learned from the fine-tuning dataset; mitigating any toxicity inherited from the base model or ensuring the chatbot inherits the base model’s safety properties [57] is a non-goal.
@@
The defender is unaware of the distribution/type or proportion of toxic language present in the training set. This is a practical and challenging setting, given the diverse ways in which toxic conversations can manifest and evolve over time [73]. The entire training pipeline (e.g., learning algorithm, hyperparameters, training infrastructure), except the training dataset, is considered trustworthy. The inference pipeline after training is also trustworthy and cannot be tampered with by the attacker.
Figure 2: Architecture of TuneShield framework.
II-C Related work
Using ML to filter toxic conversations in training data.    An obvious approach to mitigate toxicity is using an ML-based toxic language classifier to filter toxic conversations. Most prior work relies on supervised learning, assuming access to a labeled dataset with a distribution similar to the target toxic language [6, 55, 80, 79, 70] e.g., BERT-based toxicity classifier [28]. However, toxicity classifiers are known to generalize poorly to toxic language distributions that differ from those in the training dataset [29], and we also observe this in our experiments (Section V-C). In our work, we do not assume knowledge of the toxic language distribution. We also show that our approach can work with imperfect toxicity classifiers, which can derail existing defense schemes. Moreover, traditional NLP-based toxicity classifiers are highly vulnerable to adversarial samples that can evade detection [83, 22, 78]. Prior work has also studied toxicity classifiers in a non-conversational setting, i.e., determining whether a given statement is toxic [28, 3, 64, 23, 77, 14, 63, 56]. However, these methods are not suitable for a conversational setting, as the toxicity of a response can depend on the context of the conversation (i.e., previous turns), e.g., a response that simply says “yes” or supports a context discussing a harmful stereotype. Recent work also proposes LLM-based approaches for toxicity detection [29, 73, 32], which we discuss in Section VI-G.
Other methods to mitigate toxicity.    Other approaches include baking in awareness of toxic language during training to steer the chatbot away from toxic responses [24, 8, 4, 39], and inference-time approaches to steer the chatbot towards non-toxic responses [24, 48, 41, 58, 51, 66]. We discuss these approaches in more detail in Section VI-G, where we compare our proposal with these existing schemes.
Measuring toxicity in Chatbots.   These works do not mitigate toxicity; instead, they focus on measuring toxicity in chatbots, e.g., via specialized queries [67, 16, 21].
III TuneShield: Design and Implementation
III-A Design challenges
The key design challenges include the following:

### v2

# Optimus: A Robust Defense Framework for Mitigating Toxicity while Fine-Tuning Conversational AI

@@
Abstract.
Customizing Large Language Models (LLMs) on untrusted datasets poses severe risks of injecting toxic behaviors. In this work, we introduce Optimus, a novel defense framework designed to mitigate fine-tuning harms while preserving conversational utility. Unlike existing defenses that rely heavily on precise toxicity detection or restrictive filtering, Optimus addresses the critical challenge of ensuring robust mitigation even when toxicity classifiers are imperfect or biased. Optimus integrates a training-free toxicity classification scheme that repurposes the safety alignment of commodity LLMs, and employs a dual-strategy alignment process combining synthetic “healing data” with Direct Preference Optimization (DPO) to efficiently steer models toward safety. Extensive evaluations demonstrate that Optimus mitigates toxicity even when relying on extremely biased classifiers (with up to 85% degradation in Recall). Optimus outperforms the state-of-the-art defense StarDSS and exhibits strong resilience against adaptive adversarial and jailbreak attacks. Our source code and datasets are available at https://github.com/secml-lab-vt/Optimus
Conversational AI, data poisoning, toxicity mitigation, AI security
journalyear: 2026copyright: ccconference: Proceedings of the Sixteenth ACM Conference on Data and Application Security and Privacy; June 23–25, 2026; Frankfurt am Main, Germanybooktitle: Proceedings of the Sixteenth ACM Conference on Data and Application Security and Privacy (CODASPY ’26), June 23–25, 2026, Frankfurt am Main, Germanydoi: 10.1145/3800506.3803509isbn: 979-8-4007-2562-3/2026/06ccs: Security and privacyccs: Computing methodologies Machine learning
1. Introduction
@@
Conversational agents, commonly known as chatbots, are ubiquitous today—found in web applications, vehicles, mobile devices, and smart home IoT devices. Recent advances in foundation models (Bommasani et al., 2021), i.e., Large Language Models (LLMs), have transformed conversational AI. Chatbots are built by customizing or fine-tuning LLMs on application-specific conversational datasets. This enables highly effective chatbots that benefit from the emergent abilities of LLMs to follow instructions, better understand the contextual nuances of input, and generate human-like responses. The demand for customization has led companies such as Amazon (Amazonbedrock, 2024), Microsoft (azureai, 2024), and OpenAI (openai-finetuning, 2024) to offer LLM customization as a service.
@@
Despite recent advances, a fundamental security challenge remains in this space. The fine-tuning dataset used to adapt the LLM can be untrustworthy and contain problematic conversations or toxic language. Prior work has rigorously studied how an attack that poisons the training dataset with toxic language can controllably inject toxicity into a chatbot (Weeks et al., 2023), i.e., the chatbot learns to produce toxic responses. This can cause real harm, particularly for vulnerable populations, including minorities (Gibney, 2024) and those with physical or mental health challenges (Sellman, 2023). This motivates our key research question: Can we fine-tune or customize an LLM on an untrusted conversational dataset while mitigating any toxicity that can be learned from the dataset and preserving conversation quality?
Figure 1. Mitigating toxicity using Optimus.
@@
Solving this problem requires tackling multiple technical challenges: (1) The defender is unaware of the toxic language distribution in the fine-tuning dataset, making it harder to build an effective toxic language classifier, which would have been the simplest defense—i.e., simply filtering out toxic samples before training. (2) LLM model architectures, their training paradigms, and associated model customization schemes (Hu et al., 2021; Dettmers et al., 2023) are constantly evolving. Any defense framework should not be tied to specific model architectures or customization schemes. (3) Mitigating toxicity while preserving conversation quality is difficult. For example, an aggressive filter can lead to false positives that can degrade conversation quality. The defender may also want to reinforce certain desirable conversational traits, e.g., prosocial behavior.
@@
We propose Optimus, a novel defense framework that can be seamlessly integrated into existing fine-tuning pipelines, as illustrated in Figure 1. Optimus aims to mitigate any toxicity that can be learned from the untrusted dataset, while maintaining conversation quality. Our key contributions are as follows:
@@
(1) Any toxicity mitigation framework requires a module to provide some signal to identify potentially toxic samples. Similar to prior work, we rely on an ML-based toxicity classifier. However, we further innovate by leveraging recent advances in LLMs, mainly instruction tuning and safety alignment mechanisms, to build a performant toxicity classifier. Our training-free toxicity classification scheme requires no labeled reference datasets. We show that safety-aligned LLMs (e.g., LLaMA-2-Chat (Touvron et al., 2023)) can be easily adapted to detect toxic conversation samples by specifically exploiting their safety properties.
@@
(2) Even if we have a performant toxicity classifier today, it is likely to degrade in performance as toxic language evolves over time. Therefore, we need mechanisms in Optimus that can work even with imperfect or biased toxicity classifiers. To address this challenge, we propose two strategies: (a) We propose to use carefully crafted synthetic conversation samples to replace potential toxic samples in the fine-tuning dataset. We call this “healing data”. Such healing data can also be used to reinforce certain desired conversational traits, e.g., prosocial behavior. (b) We propose a model alignment mechanism based on Direct Preference Optimization (DPO) (Rafailov et al., 2023) that nudges the chatbot to learn non-toxic responses during training. This involves creating a preference dataset using our healing data samples and identified toxic samples (using our toxicity classifiers). This process is more efficient and has reduced overhead compared to methods like Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022), which require significant effort to curate human preference data. A key strength of using DPO is that even with a limited/biased set of toxic samples detected by a biased classifier, DPO can still steer the model towards the preference data distribution and generalize. Optimus is the first end-to-end scheme that can effectively mitigate toxicity, even while using an imperfect or biased toxicity classifier.
(3) We rigorously evaluate Optimus to understand its efficacy:
  • We perform a comprehensive evaluation of our toxicity classification scheme, providing insight into how well safety-aligned LLMs can perform toxicity classification. Our toxicity classifiers outperform state-of-the-art industry APIs from OpenAI (ope, 2023) and Perspective (per, 2023) by up to 28.4%.
  • We evaluate the effectiveness of individual components of Optimus, including the impact of our healing data and the DPO-based model alignment process. Optimus successfully mitigates toxicity even in the most challenging settings where we use a biased (i.e., imperfect) toxicity classifier with up to 85% degradation in Recall for a toxic sub-category.
  • Optimus outperforms the state-of-the-art defense, StarDSS (Peng et al., 2025). StarDSS’s strong reliance on an off-the-shelf toxicity classifier to dynamically shape safety during fine-tuning suffers from severe limitations when the classifier fails to accurately identify toxic samples (i.e., an imperfect classifier). This further highlights Optimus’s strength in providing safety even while using imperfect toxicity classifiers.
  • We assess the adversarial robustness of Optimus against adaptive attackers. This includes adversarial and jailbreak attacks designed to evade various stages of our pipeline. Adversarial attacks, specifically PromptAttack (Xu et al., 2023), target the toxicity classification stage. We investigate various jailbreak attacks that seek to bypass the safety alignment of LLMs involved in both classification and the generation of healing data. We also investigate Optimus’s ability to mitigate adaptive attacks in a Dialog-based learning setting (Weeks et al., 2023), where toxicity is injected into a model after deployment. Optimus even without explicit defense measures against such adaptive attacks, demonstrates resilience.
2. Background and Threat Model
2.1. Chatbots and toxicity
@@
LLMs.   We study chatbots built on foundation models (LLMs) (Bommasani et al., 2021). LLMs are pre-trained on massive text corpora and provide an excellent platform for further adaptation for a variety of downstream tasks. Modern LLMs are typically instruction-tuned to enhance instruction-following capabilities (Chung et al., 2022) and safety-aligned using RLHF (Ouyang et al., 2022) to ensure helpful and harmless behavior. A notable example is LLaMA-2-Chat (Touvron et al., 2023).
@@
Customizing LLMs to build chatbots.   LLMs are not typically designed for any particular task and therefore require further adaptation or customization for downstream use. Foundation model-customization is driving the development of generative AI applications for various use cases (Bommasani et al., 2021). Our focus is on chatbots designed for various applications, e.g., socializing, answering queries on specific or broad topics.
@@
A chatbot is built by customizing a publicly available foundation model (LLM) on a conversational dataset tailored to a specific application, e.g., customer support. A popular method for model customization is full model fine-tuning, which updates all the LLM’s parameters on a new dataset (Lewis et al., 2020). Given a conversational dataset organized in the form of context-response pairs $(X_{i},Y_{i})$, where context $X_{i}$ represents the history of previous $k$ turns of utterances $\{x_{1},x_{2},...,x_{k}\}$, and response $Y_{i}$ is the subsequent turn $x_{k+1}$. Based on the underlying LLM’s design, it is fine-tuned with an Autoregressive (Brown et al., 2020) or a Seq2Seq (Raffel et al., 2020) objective to generate a response $Y_{i}$ given the context $X_{i}$.
@@
The increasing demand for customized models spurred the development of efficient fine-tuning methods. This includes parameter-efficient fine-tuning (PEFT) techniques such as Low-Rank Adaptation (LoRA) (Hu et al., 2021). While full model fine-tuning incurs significant computational overhead with LLMs, the widely adopted LoRA technique reduces this overhead using lower-rank trainable decomposition matrices to approximate gradient updates while keeping the original model parameters frozen.
@@
Toxicity in chatbots.    Similar to prior work (Weeks et al., 2023), we consider a chatbot as toxic if its responses, interpreted within the context of the recent conversation (i.e., the last few turns), have the potential to cause harm or distress to users. Toxic language refers to any language that is harmful to a user. Rather than strictly adhering to a single, specific definition of toxicity, our analysis encompasses various types of harmful language. In Section 4, we evaluate Optimus on various datasets that cover various forms of toxicity.
@@
2.2. Defender and threat model
@@
The adversary aims to inject toxicity into a chatbot by poisoning the training (fine-tuning) dataset, i.e., by injecting toxic conversations. Prior work has examined the practicality of such toxicity injection attacks (Weeks et al., 2023), highlighting the need for mitigation schemes. Only a small fraction of toxic (poisoned) samples, e.g., 1% of the training set, is sufficient to inject toxicity into a chatbot (Weeks et al., 2023). In Section 7, we consider an adversary who is aware of our defense, and poisons the training dataset with adversarial samples to bypass our defense.
@@
Training data can be poisoned in several ways: (1) The adversary creates a poisoned conversation dataset and shares it on online repositories such as HuggingFace (hf_, 2024). (2) The adversary injects toxic conversations into forums/portals (as a user) from which they are then scraped to prepare training data. Prior work has demonstrated the feasibility of such data poisoning attacks on Wikipedia to manipulate generative AI (Carlini et al., 2024). (3) The chatbot developer sources the training data from an untrusted third party, who poisons the dataset. (4) An adversary can poison training data during Dialog-based Learning (DBL), where recent user interactions are used to update a chatbot over time (Weeks et al., 2023). The adversary masquerades as normal users to inject toxic conversations during DBL. (5) Lastly, poisoning can be incidental, i.e., without an adversary, e.g., training data is collected from forums such as Reddit that are known to have toxic conversations.
@@
We propose Optimus, a defense framework to mitigate any toxicity learned during customization or fine-tuning (to create a chatbot), while preserving conversational quality. Optimus integrates with a chatbot’s fine-tuning pipeline, and assumes that the fine-tuning (training) data is untrusted and may contain toxic conversations. A chatbot trained using Optimus on a dataset poisoned with toxic samples should exhibit similar or minimal degradation of conversation quality, compared to a model trained on the same dataset without the toxic samples. The defender trusts the base model that is used for fine-tuning. Note that our goal is solely to mitigate toxicity learned from the fine-tuning dataset; mitigating any toxicity inherited from the base model or ensuring the chatbot inherits the base model’s safety properties (Qi et al., 2023) is a non-goal.
@@
The defender is unaware of the distribution/type or proportion of toxic language present in the training set. This is a practical and challenging setting, given the diverse ways in which toxic conversations can manifest and evolve over time (Vishwamitra et al., 2024). The entire training pipeline (e.g., learning algorithm, hyperparameters, training infrastructure), except the training dataset, is considered trustworthy. The inference pipeline after training is also trustworthy and cannot be tampered with by the attacker.
3. Optimus: Defense Framework
3.1. Design Goals
We have four key design goals: